# Singular value asymptotics on compact smooth Riemannian manifolds

### v1

@@ Abstract.
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,S(1+\Delta_{G})^{-\frac{d}{2p}})=c_{d,p}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)},\quad 0<p<\infty,$
where $c_{d,p}$ is a constant, $e^{-q_{G}}$ is the canonical weight on $X$, and $d\lambda$ is the Liouville measure on the cotangent bundle $T^{\ast}X$.
@@ 1.1. Motivation on Weyl’s law
A powerful modern reformulation of Weyl’s law is expressed in terms of singular value asymptotics. Let $X$ be a $d$-dimensional compact smooth Riemannian manifold without boundary, and let $T$ be a pseudo-differential operator of order p-p on $X$ with $p>0$. In [14, Lemma 4.5], Grubb states (she referred to [1] without proof) that
@@ 1.3. Advantages and obstacles
In contrast to (1.1), the advantage of Theorem 1.2 is threefold: (a) it works on compact smooth Riemannian manifolds, including those with boundary; (b) it encompasses a strictly larger class of operators; (c) it is situated within the category of $C^{\ast}$-algebras (i.e. non-commutative topological spaces), not in a category of classical pseudodifferential operators which does not have a natural counterpart in non-commutative geometry. In addition, Theorem 1.2 covers a counterpart of Theorem A on classical tori [36] and provides Connes’ noncommutative integral for positive operators in $\Pi_{X}$.
@@ 2.2. Pseudo-differential operator
$|D_{x}^{\alpha}D_{\xi}^{\beta}a(x,\xi)|\leq c_{\alpha,\beta}(1+|\xi|_{2})^{\frac{m-|\beta|_{1}}{2}},\quad\alpha,\beta\in\mathbb{Z}_{+}^{d}.$
Here $|\beta|_{1}=\sum_{k=1}^{d}|\beta_{k}|$ and $|\xi|_{2}=(\sum_{k=1}^{d}|\xi_{k}|^{2})^{\frac{1}{2}}$. The elements in $\Psi^{m}(\mathbb{R}^{d})$ are called pseudo-differential operators of order $m$. The key properties are that
$\Psi^{m_{1}}(\mathbb{R}^{d})\cdot\Psi^{m_{2}}(\mathbb{R}^{d})\subset\Psi^{m_{1}+m_{2}}(\mathbb{R}^{d}),\quad[\Psi^{m_{1}}(\mathbb{R}^{d}),\Psi^{m_{2}}(\mathbb{R}^{d})]\subset\Psi^{m_{1}+m_{2}-1}(\mathbb{R}^{d})$
@@ 2.3. Sobolev space
is said to be elliptic if $\sum_{k,l=1}^{d}a_{kl}(x)\xi_{k}\xi_{l}\neq 0$ for all $\xi_{k}\neq 0$ ($1\leq k\leq d$). The Sobolev spaces $W^{2,2}(\mathbb{R}^{d})$ is the domain of the positive operator $\Delta$. Moreover, we crucially use the following fundamental result related to elliptic differential operator which is organised from [44].
@@ 3. Compactness of commutator on Euclidean space
where $R_{j}:=\partial_{j}\Delta^{-\frac{1}{2}}$ is the Riesz transform on $\mathbb{R}^{d}$ for $j=1,\cdots,d$.
@@ 3.1. Double operator integral
$(1+P)^{-\frac{s}{4}}[M_{f}D_{l},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
@@
Write $\phi=\phi\theta$ with $\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$. We have $f\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$. Write
$[M_{f}D_{l},M_{\phi}R_{j}M_{\phi}]=[M_{f\theta}D_{l},M_{\phi}R_{j}M_{\phi}].$
Clearly, $M_{f\theta}D_{l}\in\Psi^{1}(\mathbb{R}^{d})$ and $M_{\phi}R_{j}M_{\phi}\in\Psi^{0}(\mathbb{R}^{d})$. We have $M_{f\theta}D_{l},M_{\phi}R_{j}M_{\phi}]\in\Psi^{0}(\mathbb{R}^{d})$ and therefor it is bounded on $L_{2}(\mathbb{R}^{d})$ by [43, Theorem 2,page 286]. Since $(1+P)^{-\frac{s}{4}}$ and $(1+P)^{\frac{s}{4}-\frac{1}{2}}$ are bounded for $0\leq s\leq 2$, the assertion follows. ∎
Lemma 3.13.
$(1+P)^{-\frac{s}{4}}[M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
Since $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, there is a function $\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi=\phi\theta$ and
$M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]=[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}].$
We have $M_{f\theta}D_{i}D_{k}\in\Psi^{2}(\mathbb{R}^{d})$ and therefore $M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]\in\Psi^{1}(\mathbb{R}^{d})$. Observe that
$M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]=M_{\theta}[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]M_{\theta}.$
Thus,
$(1+\Delta)^{-\frac{s}{4}}[M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+\Delta)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+\Delta)^{-\frac{s}{4}}M_{\theta}\cdot[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]\cdot M_{\theta}(1+\Delta)^{\frac{s}{4}-\frac{1}{2}}$
$\in\Psi^{-\frac{s}{2}}(\mathbb{R}^{d})\cdot\Psi^{1}(\mathbb{R}^{d})\cdot\Psi^{\frac{s}{2}-1}(\mathbb{R}^{d})$
$\subset\Psi^{0}(\mathbb{R}^{d}).$
Combining this fact with the boundedness of operators $(1+P)^{-\frac{s}{4}}(1+\Delta)^{\frac{s}{4}}$ and $(1+\Delta)^{\frac{s}{4}-\frac{1}{2}}(1+P)^{\frac{s}{4}-\frac{1}{2}}$, [43, Theorem 2,page 286] gives the assertion. ∎
@@
It suffices to prove that the factor on the right hand side is finite. Write
$P=M_{f_{0}}+\sum_{l=1}^{d}M_{f_{l}}D_{l}+\sum_{i,k=1}^{d}M_{f_{ik}}D_{i}D_{k},$
where functions $\{f_{ik}\}_{i,k=1}^{d},$ $\{f_{l}\}_{l=1}^{d}$ and $f_{0}$ are smooth and constant outside some ball on $\mathbb{R}^{d}$. Thus,
$(1+P)^{-\frac{s}{4}}[P,M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+P)^{-\frac{s}{4}}[M_{f_{0}},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}+$
$+\sum_{l=1}^{d}(1+P)^{-\frac{s}{4}}[M_{f_{l}}D_{l},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}+$
$+\sum_{i,k=1}^{d}(1+P)^{-\frac{s}{4}}[M_{f_{ik}}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}.$
By Lemma 3.12, the second factor on the right hand side is bounded. By Lemma 3.13, the third factor on the right hand side is bounded. Since $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, write $\phi=\phi\psi$ with $\psi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By $f_{0}\psi\in C_{c}^{\infty}(\mathbb{R}^{d})$, we infer from [9] that
$M_{f_{0}},M_{\phi}R_{j}M_{\phi}]=M_{\phi}[M_{f_{0}\psi},R_{j}]M_{\phi}\in B(L_{2}(\mathbb{R}^{d}))$
and therefore, for $s\in(0,2)$,
$(1+P)^{-\frac{s}{4}}[M_{f_{0}},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
This is the estimate for the first factor and the proof is complete. ∎
@@ 5.2. Application of abstract estimate
Let us turn to the density of differentiable elements. Notably, the algebra $\{g(\frac{\nabla}{\sqrt{\Delta}}):g\in C(\mathbb{S}^{d-1})\}$ can also be generated by the Riesz transforms $R_{j}=\partial_{j}\Delta^{-\frac{1}{2}}$, $j=1,\cdots,d$. We have below results.
@@
Let $P$ be a second order differential operator with smooth compactly supported coefficients. Let us firstly claim that $\{R_{j}\}_{j=1}^{d}$ are twice differentiable. Clearly, $P\in\Psi^{2}(\mathbb{R}^{d})$ and $R_{j}\in\Psi^{0}(\mathbb{R}^{d})$ for $j=1,\cdots,d$. By gradation of pseudo-differentiable operators,
$P,R_{j}](1+\Delta)^{-\frac{1}{2}},\quad[P,[P,R_{j}]](1+\Delta)^{-1}\in\Psi^{0}(\mathbb{R}^{d})$
where we use the fact $(1+\Delta)^{-\frac{1}{2}}\in\Psi^{-1}(\mathbb{R}^{d})$ and $(1+\Delta)^{-1}\in\Psi^{-2}(\mathbb{R}^{d})$. The boundedness on $L_{2}(\mathbb{R}^{d})$ of these two commutators follows from [43, Theorem 2,page 286] and then the claim follows.
Similarly, the operator $M_{f}$ with $f\in C_{c}^{\infty}(\mathbb{R}^{d})$ is twice differentiable since $M_{f}\in\Psi^{0}(\mathbb{R}^{d})$.
Let $S\in\Pi_{\mathbb{R}^{d}}$. By the definition of $\Pi_{\mathbb{R}^{d}},$ we can find a sequence $\{S_{n}\}_{n\geq 0}$ such that each $S_{n}$ belongs to $\ast$-algebra generated by $\{M_{f}:f\in C_{0}(\mathbb{R}^{d})\}$ and by $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n}\to S$ in the uniform norm. Since $C_{c}^{\infty}(\mathbb{R}^{d})$ is dense in $C_{0}(\mathbb{R}^{d})$ in the uniform norm, there is a sub-sequence $\{S_{n_{k}}\}_{k\geq 0}$ generated by $\{M_{f}:f\in C_{c}^{\infty}(\mathbb{R}^{d})\}$ and $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n_{k}}\to S$ in the uniform norm. By arguments in the preceding two paragraphs, $S_{n_{k}}$ is twice differentiable. The proof is complete. ∎
@@
in the uniform norm. Replacing $S_{n}$ with $M_{\phi_{i}}S_{n}M_{\phi_{i}}$ if needed, we may assume without loss of generality that each $S_{n}$ is compactly supported in $h_{i}(\mathcal{U}_{i}).$
The operator $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\in\Pi_{X}$ is twice differentiable for every $n\geq 1.$ Clearly, $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\to S$ in the uniform norm. This completes the proof. ∎
@@
The approximation for summands in the first sum on the right hand side are given by Lemma 5.8. Observe that the summands in the second sum on the right hand side are compact operators. Since integral operators with smooth integral kernels are differentiable, compact and dense among compact operators in the uniform norm, the assertion follows. ∎

### v2

@@ Abstract.
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,S(1+\Delta_{G})^{-\frac{d}{2p}})=(2\pi\sqrt[d]{d})^{-\frac{1}{p}}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)},$
where $0<p<\infty$, $e^{-q_{G}}$ is the canonical weight on $X$, and $d\lambda$ is the Liouville measure on the cotangent bundle $T^{\ast}X$.
@@ 1.1. Motivation on Weyl’s law
A powerful modern reformulation of Weyl’s law is expressed in terms of singular value asymptotics. Let $X$ be a $d$-dimensional closed Riemannian manifold, and let $T$ be a pseudo-differential operator of order p-p on $X$ with $p>0$. In [14, Lemma 4.5], Grubb states (she referred to [1] without proof) that
@@ 1.3. Advantages and obstacles
In contrast to (1.1), the advantage of Theorem 1.2 is twofold: (a) it encompasses a strictly larger class of operators; (b) it is situated within the category of $C^{\ast}$-algebras (i.e. non-commutative topological spaces), not in a category of classical pseudodifferential operators which does not have a natural counterpart in non-commutative geometry. In addition, Theorem 1.2 covers a counterpart of Theorem A on classical tori [36] and provides a link with [22] Connes’ noncommutative integral for positive operators in $\Pi_{X}$.
@@ 2.2. Pseudo-differential operator
$|D_{x}^{\alpha}D_{\xi}^{\beta}a(x,\xi)|\leq c_{\alpha,\beta}(1+|\xi|)^{m-|\beta|_{1}},\quad\alpha,\beta\in\mathbb{Z}_{+}^{d}.$
Here $|\beta|_{1}=\sum_{k=1}^{d}|\beta_{k}|$ and $|\xi|=(\sum_{k=1}^{d}|\xi_{k}|^{2})^{\frac{1}{2}}$. The elements in $\Psi^{m}(\mathbb{R}^{d})$ are called pseudo-differential operators of order $m$. The key properties are that
(2.1) $\Psi^{m_{1}}(\mathbb{R}^{d})\cdot\Psi^{m_{2}}(\mathbb{R}^{d})\subset\Psi^{m_{1}+m_{2}}(\mathbb{R}^{d}),\quad[\Psi^{m_{1}}(\mathbb{R}^{d}),\Psi^{m_{2}}(\mathbb{R}^{d})]\subset\Psi^{m_{1}+m_{2}-1}(\mathbb{R}^{d})$
@@ 2.3. Sobolev space
is said to be elliptic if the matrix $(a_{kl}(x))_{k,l=1}^{d}$ is positive definite for all $x\in\mathbb{R}^{d}$. The Sobolev spaces $W^{2,2}(\mathbb{R}^{d})$ is the domain of the positive operator $\Delta$. Moreover, we crucially use the following fundamental result related to elliptic differential operator which is organised from [44].
@@ 3. Compactness of commutator on Euclidean space
where $R_{j}:=D_{j}\Delta^{-\frac{1}{2}}$ is the Riesz transform on $\mathbb{R}^{d}$ for $j=1,\cdots,d$.
@@ 3.1. Double operator integral
$(1+P)^{-\frac{s}{4}}[P,M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
@@
Denote
$A_{1}=M_{\phi}D_{j}(1+\Delta)^{-\frac{1}{2}}M_{\phi}\quad\mbox{and}\quad A_{2}=M_{\phi}D_{j}(\Delta^{-\frac{1}{2}}-(1+\Delta)^{-\frac{1}{2}})M_{\phi}.$
Write
$[P,M_{\phi}R_{j}M_{\phi}]=[P,A_{1}]+[P,A_{2}].$
By the assumption on $P$, we have $P\in\Psi^{2}(\mathbb{R}^{d})$. Since $D_{j}(1+\Delta)^{-\frac{1}{2}}\in\Psi^{0}(\mathbb{R}^{d})$ and $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, it follows from (2.1) that
$(1+\Delta)^{-\frac{s}{4}}[P,A_{1}](1+\Delta)^{\frac{s}{4}-\frac{1}{2}}\in\Psi^{0}(\mathbb{R}^{d}),$
where we use the fact that
$(1+\Delta)^{\frac{t}{2}}\in\Psi^{t}(\mathbb{R}^{d}),\quad\forall\;t\in\mathbb{R}.$
By [43, Theorem 1,page 235], $(1+\Delta)^{-\frac{s}{4}}[P,A_{1}](1+\Delta)^{\frac{s}{4}-\frac{1}{2}}$ is bounded on $L_{2}(\mathbb{R}^{d})$. Therefore, the operator
$(1+P)^{-\frac{s}{4}}[P,A_{1}](1+P)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+P)^{-\frac{s}{4}}(1+\Delta)^{\frac{s}{4}}\cdot(1+\Delta)^{-\frac{s}{4}}[P,A_{1}](1+\Delta)^{\frac{s}{4}-\frac{1}{2}}\cdot(1+\Delta)^{\frac{1}{2}-\frac{s}{4}}(1+P)^{\frac{s}{4}-\frac{1}{2}}$
is bounded on $L_{2}(\mathbb{R}^{d})$ because the boundedness for the first and third factors on the right hand side of the equality follows from Theorem 2.1, for the second one follows by previous argument. Turn to $P,A_{2}]$. Observe that
$(1+P)^{-\frac{s}{4}}PA_{2}(1+P)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+P)^{-\frac{s}{4}}(1+\Delta)^{\frac{s}{4}}\cdot(1+\Delta)^{-\frac{s}{4}}PM_{\phi}(1+\Delta)^{\frac{s}{4}-\frac{3}{2}}\cdot\frac{R_{j}\sqrt{1+\Delta}}{\sqrt{1+\Delta}+\sqrt{\Delta}}\cdot$
$\quad\cdot(1+\Delta)^{\frac{1}{2}-\frac{s}{4}}M_{\phi}(1+\Delta)^{\frac{s}{4}-\frac{1}{2}}\cdot(1+\Delta)^{\frac{1}{2}-\frac{s}{4}}(1+P)^{\frac{s}{4}-\frac{1}{2}}.$
On right hand side of the equality, the first and fifth factors are bounded by Theorem 2.1, the second and fourth factors belong to $\Psi^{0}(\mathbb{R}^{d})$ and therefore are bounded, the third factor is obviously bounded. Thus, $(1+P)^{-\frac{s}{4}}PA_{2}(1+P)^{\frac{s}{4}-\frac{1}{2}}$ is a bounded operator. Similarly, $(1+P)^{-\frac{s}{4}}A_{2}P(1+P)^{\frac{s}{4}-\frac{1}{2}}$ is also bounded. In a sum, the operator $(1+P)^{-\frac{s}{4}}[P,A_{2}](1+P)^{\frac{s}{4}-\frac{1}{2}}$ is bounded on $L_{2}(\mathbb{R}^{d})$. The proof is complete. ∎
@@
By Lemma 3.12, the second factor on the inequality is finite. The proof is complete. ∎
@@ 5.2. Application of abstract estimate
Let us turn to the density of differentiable elements. Notably, the algebra $\{g(\frac{\nabla}{\sqrt{\Delta}}):g\in C(\mathbb{S}^{d-1})\}$ can also be generated by the Riesz transforms $R_{j}=D_{j}\Delta^{-\frac{1}{2}}$, $j=1,\cdots,d$. We have below results.
@@
Let $P$ be a second order differential operator with smooth compactly supported coefficients. Let $f\in C_{c}^{\infty}(\mathbb{R}^{d})$. By (2.1),
$P,M_{f}](1+\Delta)^{-\frac{1}{2}},\quad[P,[P,M_{f}]](1+\Delta)^{-1}\in\Psi^{0}(\mathbb{R}^{d}).$
Thus, the commutators on the left hand side are bounded, and therefore $M_{f}$ is twice differentiable.
Let’s next show that Riesz transforms $\{R_{j}\}_{j=1}^{d}$ are twice differentiable. It suffices to show
$P,R_{j}](1+\Delta)^{-\frac{1}{2}},\quad[P,[P,R_{j}]](1+\Delta)^{-1}\in B(L_{2}(\mathbb{R}^{d})).$
The boundedness for the first operator can be obtained by adapting the proof of Lemma 3.12. For the second one, write
$P,[P,R_{j}]]=[P,[P,D_{j}(1+\Delta)^{-\frac{1}{2}}]]+[P,[P,D_{j}(\Delta^{-\frac{1}{2}}-(1+\Delta)^{-\frac{1}{2}})]].$
Clearly, $P\in\Psi^{2}(\mathbb{R}^{d})$ as $P$ has smooth compactly supported coefficients. By (2.1),
$P,[P,D_{j}(1+\Delta)^{-\frac{1}{2}}]](1+\Delta)^{-1}\in\Psi^{0}(\mathbb{R}^{d}).$
The boundedness of this operator follows from [43, Theorem 1,page 235]. Observe that
$A:=D_{j}(\Delta^{-\frac{1}{2}}-(1+\Delta)^{-\frac{1}{2}})=(1+\Delta)^{-1}\cdot\frac{R_{j}\sqrt{1+\Delta}}{\sqrt{1+\Delta}+\sqrt{\Delta}},$
and
$P,[P,A]]=P^{2}A-2PAP+AP^{2}.$
Write
$P^{2}A(1+\Delta)^{-1}=P^{2}(1+\Delta)^{-2}\cdot\frac{R_{j}\sqrt{1+\Delta}}{\sqrt{1+\Delta}+\sqrt{\Delta}}.$
On the right hand side of the equality, the first factor belongs to $\Psi^{0}(\mathbb{R}^{d})$ due to $P^{2}\in\Psi^{4}(\mathbb{R}^{d})$ and (2.1), and therefore is bounded on $L_{2}(\mathbb{R}^{d})$. The boundedness for the second one is obvious. Similarly, the operators
$PAP(1+\Delta)^{-1}=P(1+\Delta)^{-1}\cdot\frac{R_{j}\sqrt{1+\Delta}}{\sqrt{1+\Delta}+\sqrt{\Delta}}\cdot P(1+\Delta)^{-1}$
and
$AP^{2}(1+\Delta)^{-1}=\frac{R_{j}\sqrt{1+\Delta}}{\sqrt{1+\Delta}+\sqrt{\Delta}}\cdot(1+\Delta)^{-1}P^{2}(1+\Delta)^{-1}$
are bounded on $L_{2}(\mathbb{R}^{d})$. In a sum, $P,[P,A]](1+\Delta)^{-1}$ is bounded on $L_{2}(\mathbb{R}^{d})$. Moreover, a polynomial $S$ generated by $\{M_{f}:f\in C_{c}^{\infty}(\mathbb{R}^{d})\}$ and $\{R_{j}\}_{j=1}^{d}$ is twice differentiable.
For $S\in\Pi_{\mathbb{R}^{d}}$, pick a sequence $\{S_{n}\}_{n\geq 0}$ such that each $S_{n}$ belongs to $\ast$-algebra generated by $\{M_{f}:f\in C_{0}(\mathbb{R}^{d})\}$ and by $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n}\to S$ in the uniform norm. Since $C_{c}^{\infty}(\mathbb{R}^{d})$ is dense in $C_{0}(\mathbb{R}^{d})$ in the uniform norm, there is a sub-sequence $\{S_{n_{k}}\}_{k\geq 0}$ generated by $\{M_{f}:f\in C_{c}^{\infty}(\mathbb{R}^{d})\}$ and $\{R_{j}\}_{j=1}^{d}$, and $S_{n_{k}}\to S$ in the uniform norm. By the previous arguments, $S_{n_{k}}$ is twice differentiable. The assertion follows. ∎
@@
in the uniform norm. Replacing $S_{n}$ with $M_{\phi_{i}}S_{n}M_{\phi_{i}}$, we may assume without loss of generality that each $S_{n}$ is compactly supported in $h_{i}(\mathcal{U}_{i}).$ In this case, $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\to S$ in the uniform norm and therefore it suffices to show that $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\in\Pi_{X}$ is twice differentiable for every $n\geq 1.$
Let $P$ be a second order differential operator with smooth coefficients on $X$. Let us prove that
(5.8) $[W_{h_{i}}^{*}S_{n}W_{h_{i}},P]:W^{1,2}(X,{\rm vol}_{G})\rightarrow L_{2}(X,{\rm vol}_{G})$
and
(5.9) $[[W_{h_{i}}^{*}S_{n}W_{h_{i}},P],P]:W^{2,2}(X,{\rm vol}_{G})\rightarrow L_{2}(X,{\rm vol}_{G})$
are bounded. We begin by considering (5.8). Denote $\psi_{i}=\phi_{i}\circ h_{i}$ and write
(5.10) $\begin{split}&[W_{h_{i}}^{*}S_{n}W_{h_{i}},P]\\ &=W_{h_{i}}^{*}S_{n}W_{h_{i}}[M_{\psi_{i}},P]+[W_{h_{i}}^{*}S_{n}W_{h_{i}},M_{\psi_{i}}PM_{\psi_{i}}]+[P,M_{\psi_{i}}]W_{h_{i}}^{*}S_{n}W_{h_{i}}.\end{split}$
The first and third factors on the right hand side of equality are bounded from $W^{1,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$ since $W_{h_{i}}^{*}S_{n}W_{h_{i}}$ is bounded on $L_{2}(X,{\rm vol}_{G})$ and $M_{\psi_{i}},P]:W^{1,2}(X,{\rm vol}_{G})\to L_{2}(X,{\rm vol}_{G})$ is bounded. For the second factor, write
$W_{h_{i}}^{*}S_{n}W_{h_{i}},M_{\psi_{i}}PM_{\psi_{i}}]=W_{h_{i}}^{*}[S_{n},P_{i}]W_{h_{i}}$
where $P_{i}=W_{h_{i}}M_{\psi_{i}}PM_{\psi_{i}}W_{h_{i}}^{*}$ is a second order differential operator with smooth coefficients compactly supported in $\Omega_{i}$. Pick $\eta_{i}\in C_{c}^{\infty}(\Omega_{i})$ such that
$S_{n},P_{i}]=M_{\eta_{i}}[S_{n},P_{i}]M_{\eta_{i}}.$
Since $S_{n}$ is twice differentiable, it follows that $S_{n},P_{i}]:W^{1,2}(\Omega_{i})\rightarrow L_{2}(\Omega_{i})$ is bounded. Clearly,
$M_{\eta_{i}}:W^{1,2}(\Omega_{i},{\rm vol}_{g_{i}})\rightarrow W^{1,2}(\Omega_{i})$
and
$M_{\eta_{i}}:L_{2}(\Omega_{i})\rightarrow L_{2}(\Omega_{i},{\rm vol}_{g_{i}})$
are bounded. In a sum, $S_{n},P_{i}]:W^{1,2}(\Omega_{i},{\rm vol}_{g_{i}})\rightarrow L_{2}(\Omega_{i},{\rm vol}_{g_{i}})$ is bounded and therefore the assertion for (5.8) holds.
Turn to investigate (5.9). Firstly,
$W_{h_{i}}^{*}S_{n}W_{h_{i}}[M_{\psi_{i}},P],P]=W_{h_{i}}^{*}S_{n}W_{h_{i}}[[M_{\psi_{i}},P],P]+[W_{h_{i}}^{*}S_{n}W_{h_{i}},P][M_{\psi_{i}},P].$
The first factor on the right hand side is bounded from $W^{2,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$ since $M_{\psi_{i}},P],P]:W^{2,2}(X,{\rm vol}_{G})\to L_{2}(X,{\rm vol}_{G})$ is bounded and $W_{h_{i}}^{*}S_{n}W_{h_{i}}$ is bounded on $L_{2}(X,{\rm vol}_{G})$. The second one is bounded from $W^{2,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$ since operator $M_{\psi_{i}},P]:W^{2,2}(X,{\rm vol}_{G})\to W^{1,2}(X,{\rm vol}_{G})$ is bounded and $W_{h_{i}}^{*}S_{n}W_{h_{i}},P]:W^{1,2}(X,{\rm vol}_{G})\to L_{2}(X,{\rm vol}_{G})$ is bounded by (5.8). Similarly,
$P,M_{\psi_{i}}]W_{h_{i}}^{*}S_{n}W_{h_{i}},P]=+[P,M_{\psi_{i}}][W_{h_{i}}^{*}S_{n}W_{h_{i}},P]+[[P,M_{\psi_{i}}],P]W_{h_{i}}^{*}S_{n}W_{h_{i}}$
is bounded from $W^{2,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$. Let $A_{i}=[W_{h_{i}}^{*}S_{n}W_{h_{i}},M_{\psi_{i}}PM_{\psi_{i}}]$. We have $A_{i}=M_{\psi_{i}}A_{i}M_{\psi_{i}}$. Moreover,
$A_{i},P]=A_{i}[M_{\psi_{i}},P]+[A_{i},M_{\psi_{i}}PM_{\psi_{i}}]+[P,M_{\psi_{i}}]A_{i}.$
The first factor is bounded from $W^{2,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$ since $M_{\psi_{i}},P]$ is bounded from $W^{2,2}(X,{\rm vol}_{G})$ to $W^{1,2}(X,{\rm vol}_{G})$ and $A_{i}$ is bounded from $W^{1,2}(X,{\rm vol}_{G})$ to $L_{2}(X,{\rm vol}_{G})$. The third factor is similar. For the second one, write
$A_{i},M_{\psi_{i}}PM_{\psi_{i}}]=W_{h_{i}}^{*}[[S_{n},P_{i}],P_{i}]W_{h_{i}}$
where $P_{i}=W_{h_{i}}M_{\psi_{i}}PM_{\psi_{i}}W_{h_{i}}^{*}$. Since $S_{n}$ is twice differentiable, it follows that $S_{n},P_{i}],P_{i}]:W^{2,2}(\Omega_{i},{\rm vol}_{g_{i}})\rightarrow L_{2}(\Omega_{i},{\rm vol}_{g_{i}})$ is bounded by adapting the arguments in the last second paragraph. Lastly, by (5.10), the assertion for (5.9) holds. This completes the proof. ∎
@@
The approximation for summands in the first sum on the right hand side are given by Lemma 5.8. Observe that the summands in the second sum on the right hand side are compact operators. Since integral operators with smooth integral kernels are twice differentiable, compact and dense among compact operators in the uniform norm, the assertion follows. ∎